If the total distance of a journey is 120 km .If one goes by 60 kmph and comes back at 40kmph what is the average speed during the journey?

Answer by richwmiller(17219)

It is not (60+40)/2=50
Why? because the times at each speed are different.
It will be close to 50 though since the difference in times is so small.
60*x=120
40*y=120
x=2 hrs
y=3 hrs
total distance over total time
total time is 5 hours
120+120=240 total distance
240/5=48 kph